### 28 Apr 2026 v2

# Thermodynamic Curvature and the Widom Ridge in Interacting Spin Systems

Abstract
We develop a geometric formulation of thermodynamic response in the classical Ising model by defining a curvature field over the control manifold spanned by inverse temperature $\beta$ and magnetic field $h$. We show that the existence of nontrivial curvature depends sensitively on the choice of control variables: while the $(J,h)$ manifold at fixed temperature is integrable and exhibits zero curvature, the $(\beta,h)$ manifold supports a finite curvature field arising from variations of the statistical ensemble. This curvature is given by a mixed derivative of the free energy and can be expressed directly as the covariance between energy and magnetization fluctuations. We evaluate the curvature field using Monte Carlo sampling and demonstrate that it develops a pronounced ridge structure extending from the critical point into the supercritical regime. This identifies the Widom line as a geometric feature of control space, corresponding to a locus of maximal thermodynamic response. More generally, the formulation provides a direct connection between geometric thermodynamics, critical phenomena, and experimentally accessible observables, and suggests that thermodynamic curvature may be probed through measurements of work performed under cyclic driving protocols.
I Introduction
Many central relations of thermodynamics admit a natural geometric interpretation. The most familiar example is the identification of mechanical work with the area enclosed by a cycle in the $(P,V)$ plane,
$W=\oint P\,dV,$ (1)
which expresses work as a line integral over a closed curve. More generally, thermodynamic relations may be formulated in the language of differential forms, where work and heat appear as integrals over trajectories in state space and the corresponding area laws arise from the associated surface integrals. This viewpoint reflects an underlying geometric structure encoded in the First Law,
dU=TdSPdV,dU=T\,dS-P\,dV, (2)
which endows the equilibrium manifold with a natural differential-geometric structure.
Our recent work has emphasized that this geometric structure may be expressed locally in terms of curvature fields defined over thermodynamic variables, such that infinitesimal work generated by small cycles is governed by a corresponding curvature density [2, 4, 3]. In this formulation, thermodynamic work is not solely a global property of cyclic processes, but rather a local geometric quantity defined over state space. While this structure is naturally expressed in terms of equilibrium variables such as $(S,T)$ or $(P,V)$, many systems of interest—particularly in statistical mechanics—are more naturally parameterized by external control variables such as temperature and field. This raises a natural question: how does the geometric structure of thermodynamics manifest in systems described by a free energy as a function of control parameters?
Geometric approaches to thermodynamics have developed along several complementary directions. In equilibrium, contact geometry provides a natural framework in which thermodynamic variables define a contact manifold and equilibrium states appear as Legendre submanifolds [12, 18, 6, 5]. Metric-based formulations due to Weinhold and Ruppeiner relate geometric structure to equilibrium fluctuations and phase behavior [24, 21]. Extensions to nonequilibrium systems, particularly within stochastic thermodynamics, describe work as a trajectory-dependent functional obeying exact fluctuation relations such as the Jarzynski equality and Crooks theorem [14, 9, 22]. These developments establish a broad geometric perspective on thermodynamics, but are typically formulated either in terms of intrinsic properties of equilibrium state space or in terms of trajectory-dependent quantities.
In contrast, the present work focuses on a curvature defined directly over the space of externally controllable variables. In this formulation, thermodynamic response is encoded in a two-form whose flux determines the work performed over cyclic processes, providing a direct connection between geometric thermodynamics, statistical mechanics, and experimentally accessible observables.
We address this question for the classical Ising model by treating the inverse temperature $\beta$ and magnetic field $h$ as coordinates on a control manifold. We show that the resulting curvature field develops pronounced structure near the critical point and becomes sharply localized along a ridge in the supercritical regime.
The concept of the Widom line has emerged as a central organizing principle for understanding thermodynamic crossover behavior in systems with critical endpoints. Originally developed in the context of supercooled water and liquid–liquid phase transitions, the Widom line is typically defined as the locus of extrema in response functions or correlation length extending into the supercritical regime [25, 11]. Experimental and simulation studies have demonstrated that this line is associated not only with thermodynamic anomalies, but also with dynamical crossovers, including the breakdown of the Stokes–Einstein relation and changes in transport properties [15, 19], establishing it as an observable feature of many-body systems [13].
From a theoretical perspective, Widom-line behavior arises from the structure of critical fluctuations. Scaling analyses show that extrema of response functions converge onto a well-defined line emanating from the critical point, reflecting the underlying universality class [17]. At the same time, symmetry considerations suggest that the structure of the supercritical region may be richer than a single crossover line [16], and extensions to modified Ising-like models and other physical systems further reinforce the universality of this behavior [7, 23].
Despite this progress, the Widom line is most often defined operationally through extrema of individual response functions. While effective, this approach obscures the fact that these extrema arise from correlated fluctuations between thermodynamic variables. Several works have explored geometric formulations in this context, including information geometry, where scalar curvature tracks phase transitions and crossover behavior [10]. In parallel, dynamical studies of Ising systems and recent lattice investigations demonstrate that Widom-like behavior can be identified through observable response beyond equilibrium susceptibilities [20, 8].
In this work, we show that the Widom line admits a natural geometric interpretation as a ridge in the curvature field defined over control space. The curvature may be expressed directly in terms of correlated fluctuations between energy and magnetization, providing a unified and observable measure of thermodynamic response.
Finally, this formulation suggests a direct experimental route to probing Widom-line behavior through controlled cyclic driving. Since the curvature is defined as the flux of geometric work over closed loops in control space, it may be accessed through measurements of work performed under small cycles in temperature and field. In this approach, the Widom ridge appears as the locus of maximal geometric response, providing an experimentally accessible signature of thermodynamic curvature in interacting systems. =
II Geometric Formulation of Thermodynamic Response
We briefly summarize the geometric formulation of thermodynamic response and adapt it to systems parameterized by external control variables.
For a system described by an energy function $U(S,V)$, the First Law,
dU=TdSPdV,dU=T\,dS-P\,dV, (3)
defines a differential structure on thermodynamic state space. Local thermodynamic response is governed by mixed derivatives of the fundamental potential, such as
$\Omega_{SV}=\frac{\partial^{2}U}{\partial S\partial V},$ (4)
which characterize the coupling between thermodynamic variables and determine the response to infinitesimal variations.
More generally, for a system parameterized by a set of externally controlled variables $\lambda=(\lambda_{1},\lambda_{2})$, the work differential may be written as a one-form,
$\delta W=A_{i}(\lambda)\,d\lambda_{i},$ (5)
with associated curvature
$\mathcal{F}_{ij}=\partial_{\lambda_{i}}A_{j}-\partial_{\lambda_{j}}A_{i}.$ (6)
The work performed over a closed cycle $C$ is then given by
$W[C]=\oint_{C}A_{i}\,d\lambda_{i}=\iint_{\Sigma_{C}}\mathcal{F}_{ij}\,d\lambda_{i}\wedge d\lambda_{j},$ (7)
identifying thermodynamic work as the flux of a curvature field defined over control-parameter space.
In this formulation, a nonzero work over a cycle arises precisely when the one-form $A_{i}(\lambda)\,d\lambda_{i}$ is non-exact, corresponding to a nonvanishing curvature $\mathcal{F}_{ij}$. The existence and structure of this curvature depend sensitively on the choice of control variables.
Geometric interpretation: work as holonomy.
The work performed over a quasistatic cycle admits a natural geometric interpretation in terms of holonomy. i.e., the accumulated change obtained by transporting the system around a closed loop in parameter space, which depends on the path taken rather than only its endpoints. Writing the work as
$W[C]=\oint_{C}\mathcal{A},$ (8)
with $\mathcal{A}=A_{i}(\lambda)\,d\lambda_{i}$, the one-form $\mathcal{A}$ plays the role of a connection over control space. The corresponding curvature two-form,
$\Omega=d\mathcal{A},$ (9)
determines the local geometric response, and by Stokes’ theorem,
$W[C]=\iint_{\Sigma_{C}}\Omega.$ (10)
In the limit of infinitesimal cycles, the curvature may be identified as the work per unit area,
$\Omega_{ij}=\lim_{A\to 0}\frac{W[C]}{A}.$ (11)
This relation provides an operational interpretation of the curvature as a local density of geometric work. In this sense, quasistatic thermodynamic work constitutes a geometric holonomy, with the curvature encoding the local structure of thermodynamic response. This perspective also suggests that thermodynamic curvature may be accessed experimentally through measurements of work performed over small cycles in control space.
In the following section, we apply this framework to the Ising model and demonstrate explicitly how the choice of control variables determines the emergence of nontrivial curvature.
Relation to geometric response and Berry curvature.
Geometric response functions analogous to Berry curvature have been identified in both quantum and classical nonequilibrium systems, where cyclic driving in parameter space leads to geometric phases and modified response properties [1]. In such settings, the curvature is associated with the breakdown of equilibrium thermodynamic relations and the emergence of geometric contributions to response.
In contrast, the present formulation arises within equilibrium statistical mechanics. Here, the curvature is derived directly from thermodynamic potentials and can be expressed in terms of equilibrium fluctuations. Rather than signaling a departure from thermodynamic structure, it provides a local description of thermodynamic response and encodes the coupling between control variables.
III Geometric Structure of the Ising Model
We consider the two-dimensional Ising model with Hamiltonian
$H(J,h)=-J\,\mathcal{B}-h\,\mathcal{M}$, (12)
where $\mathcal{B}=\sum_{\langle ij\rangle}s_{i}s_{j}$ and $\mathcal{M}=\sum_{i}s_{i}$. The equilibrium state is given by the Gibbs distribution $\rho\propto e^{-\beta H}$, with partition function
$Z(\beta,h)=\sum_{\{s_{i}\}}e^{-\beta H},$ (13)
and free energy
$F(\beta,h)=-\frac{1}{\beta}\ln Z(\beta,h).$ (14)
The thermodynamic response depends on how the system is parameterized in control space. We therefore distinguish two natural manifolds: the coupling–field manifold $(J,h)$ at fixed $\beta$, and the thermodynamic control manifold $(\beta,h)$ at fixed $J$.
III.1 Flat and Curved Control Manifolds: $(J,h)$ vs $(\beta,h)$
The emergence of a nontrivial curvature field depends sensitively on the choice of control variables.
The $(J,h)$ manifold at fixed $\beta$.
At fixed inverse temperature, the free energy defines an exact differential,
$dF=-\langle\mathcal{B}\rangle\,dJ-\langle\mathcal{M}\rangle\,dh,$ (15)
so that the associated one-form
$\mathcal{A}_{Jh}=-\langle\mathcal{B}\rangle\,dJ-\langle\mathcal{M}\rangle\,dh$ (16)
satisfies $\mathcal{A}_{Jh}=dF$. The curvature therefore vanishes identically,
$\Omega_{Jh}=d\mathcal{A}_{Jh}=0.$ (17)
Thus, the $(J,h)$ manifold is globally integrable: no geometric work can be generated.
The $(\beta,h)$ manifold at fixed $J$.
When $\beta$ is treated as a control parameter, the work associated with magnetic cycles is determined by the magnetization,
$\mathcal{A}=-M(\beta,h)\,dh.$ (18)
The corresponding curvature two-form is
$\Omega=d\mathcal{A}=\Omega_{\beta h}\,d\beta\wedge dh,$ (19)
with
$\Omega_{\beta h}=-\left(\frac{\partial M}{\partial\beta}\right)_{h}=-\frac{\partial^{2}F}{\partial\beta\partial h}.$ (20)
Using standard identities for canonical averages, the mixed derivative may be expressed in terms of equilibrium fluctuations. For an observable $O$,
$\frac{\partial}{\partial\beta}\langle O\rangle=-\left(\langle OH\rangle-\langle O\rangle\langle H\rangle\right).$ (21)
Applying this identity to the magnetization $M=\langle\mathcal{M}\rangle$ yields
$\Omega_{\beta h}=\langle\mathcal{M}H\rangle-\langle\mathcal{M}\rangle\langle H\rangle.$ (22)
Introducing the intensive variables $m=\mathcal{M}/N$ and $e=H/N$, we obtain
$\Omega_{\beta h}=-N\left(\langle me\rangle-\langle m\rangle\langle e\rangle\right).$ (23)
This result shows that the curvature is proportional to the covariance between energy and magnetization fluctuations. The geometric response is therefore governed by correlated fluctuations of the conjugate observables and is enhanced in regions where these fluctuations become strongly coupled, such as near the critical point.
Origin of geometric response.
The distinction between these manifolds reflects the role of $\beta$ in defining the statistical ensemble. Variations in $(J,h)$ at fixed $\beta$ deform the Hamiltonian within a fixed ensemble, leading to an exact differential structure. In contrast, variation of $\beta$ changes the Boltzmann weights of all configurations, introducing a genuinely thermodynamic degree of freedom. The resulting response is therefore non-integrable and supports a nontrivial curvature field.
Widom ridge as a geometric feature.
Above the critical temperature, the Ising model exhibits no true phase transition, but retains a well-defined line of enhanced response in the $(\beta,h)$ plane. This line is typically identified as the locus of extrema in response functions.
Within the present framework, this structure follows directly from the curvature field. From Eq. (23), $\Omega_{\beta h}(\beta,h)$ measures the covariance between energy and magnetization fluctuations, and thus the strength of their statistical coupling. We therefore define the Widom ridge as the locus
$\beta^{*}(h)=\arg\min_{\beta}\,\Omega_{\beta h}(\beta,h),$ (24)
corresponding to maximal curvature magnitude at fixed field.
From this perspective, the Widom line acquires a natural geometric interpretation as a ridge in the curvature field. As the critical point is approached, the curvature becomes increasingly localized, and the ridge sharpens into the singular structure associated with the phase transition.
This formulation also suggests a direct route to experimental identification. Since $\Omega_{\beta h}$ is proportional to the covariance between energy and magnetization fluctuations, it may be accessed through measurements of correlated response functions. In this sense, the Widom ridge corresponds to the locus of maximal cross-response, providing a direct and observable signature of thermodynamic curvature.
Figure 1: Curvature field $\phi_{\beta}(\beta,h)$ for the two-dimensional Ising model obtained from Monte Carlo sampling. The solid line denotes the Widom ridge, defined as the locus $\beta^{*}(h)=\arg\min_{\beta}\phi_{\beta}(\beta,h)$. Gray points indicate the sampling locations used in the numerical evaluation.
IV Numerical Evaluation of the Curvature Field
To evaluate the curvature field $\Omega_{\beta h}$ for the Ising model, we employ equilibrium Monte Carlo sampling of the canonical ensemble. As shown in the previous section, the curvature component may be expressed as
$\Omega_{\beta h}=-N\left(\langle me\rangle-\langle m\rangle\langle e\rangle\right),$ (25)
which provides a direct estimator in terms of equilibrium fluctuations of the intensive magnetization $m$ and energy density $e$.
IV.1 Simulation protocol
We consider the two-dimensional square-lattice Ising model with periodic boundary conditions and linear system size $L$, such that $N=L^{2}$. For each point $(\beta,h)$ in control space, equilibrium configurations are generated using a standard Metropolis algorithm. After an initial thermalization period, expectation values $\langle m\rangle$, $\langle e\rangle$, and $\langle me\rangle$ are computed via time averaging over Monte Carlo sweeps.
To resolve the structure of the curvature field, sampling is performed over a nonuniform grid in $(\beta,h)$, with increased density in regions where the curvature varies rapidly. In particular, additional sampling points are concentrated along the Widom ridge, as identified iteratively from preliminary simulations.
IV.2 Curvature field and Widom ridge
The resulting curvature field $\phi_{\beta}(\beta,h)\equiv\Omega_{\beta h}$ is shown in Fig. 1. The field is evaluated using the fluctuation-based estimator $\phi_{\beta}=-N\left(\langle me\rangle-\langle m\rangle\langle e\rangle\right)$, where $m$ and $e$ are the intensive magnetization and energy densities. The color scale indicates the magnitude and sign of the curvature, with a pronounced negative ridge extending from the critical point at $(\beta_{c},0)$ into the supercritical regime. The solid line denotes the Widom ridge, defined as the locus $\beta^{*}(h)=\arg\min_{\beta}\phi_{\beta}(\beta,h)$. The ridge closely follows the region of maximal curvature magnitude, demonstrating that the Widom line corresponds to a geometric feature of the control manifold. As $h\to 0$, the curvature becomes increasingly localized and the ridge sharpens, reflecting the enhancement of correlated energy–magnetization fluctuations near criticality. The gray points indicate the sampling locations used in the Monte Carlo evaluation, with increased density in the vicinity of the ridge to resolve the rapidly varying curvature. This figure provides direct numerical evidence that thermodynamic crossover behavior in the Ising model is governed by the structure of the curvature field.
Several key features are immediately apparent:
  • First, the curvature field exhibits a pronounced negative ridge extending from the critical point at $(\beta_{c},h=0)$ into the supercritical regime. This ridge corresponds to the region where the covariance between energy and magnetization fluctuations is maximized in magnitude, indicating strong coupling between these observables.
  • Second, the curvature becomes increasingly localized as $h\to 0$, reflecting the divergence of thermodynamic response near the critical point. Away from this region, the curvature field varies smoothly and approaches small values, indicating weak coupling between energy and magnetization fluctuations.
The Widom ridge, defined as the locus of minimal $\Omega_{\beta h}$ at fixed $h$, is overlaid in Fig. 1. The ridge closely follows the region of maximal curvature magnitude, confirming the geometric interpretation developed above. In particular, the numerical results demonstrate that the Widom line may be identified as a ridge in the curvature field, rather than solely as an extremum of a specific response function.
These results provide direct numerical evidence that the curvature field $\Omega_{\beta h}$ captures the essential thermodynamic structure of the Ising model in the supercritical regime. The identification of the Widom ridge as a geometric feature of the control manifold is supported by the strong localization of curvature and its direct connection to correlated fluctuations. Moreover, the fluctuation-based estimator in Eq. (25) allows the curvature field to be computed without recourse to numerical differentiation, providing a stable and efficient route for evaluating geometric response in interacting systems.
V Discussion
The geometric formulation developed here provides a new perspective on thermodynamic response in statistical systems. In the Ising model, the curvature field defined over $(\beta,h)$ captures the coupling between energy and magnetization fluctuations and reveals a rich geometric structure near criticality.
The identification of the Widom ridge as a locus of maximal curvature provides a direct geometric interpretation of supercritical response, linking classical thermodynamics, statistical mechanics, and geometric structure.
More broadly, this framework suggests that phase transitions and crossover phenomena may be understood in terms of curvature fields defined over control space, opening the door to geometric classification of thermodynamic behavior in complex systems.
More generally, the geometric formulation suggests that thermodynamic curvature fields can be reconstructed experimentally through local probing of control space. By performing a sequence of small, oriented cycles in $(\beta,h)$ and measuring the corresponding work, one may infer the local curvature via Stokes’ theorem, thereby mapping the geometric structure of the thermodynamic manifold. This approach is closely aligned with recent efforts to extract geometric and topological response functions in driven quantum and classical systems.
Such measurements may be feasible in a range of experimental platforms, including two-dimensional magnetic materials, cold-atom simulators of Ising-like Hamiltonians, and nanoscale systems where both temperature and field can be modulated dynamically. In these settings, the Widom ridge would be identified as a region of enhanced geometric response, providing a direct and operational probe of crossover behavior beyond conventional susceptibility-based definitions.
Data Availability Statement
All data generated or analyzed during this study are included in this manuscript.
Acknowledgements.
This work at the University of Houston was supported by the National Science Foundation under CHE-2404788 and the Robert A. Welch Foundation (E-1337).
Author Contributions
The author developed the theoretical framework, performed all derivations, and carried out the analysis presented in this work. Generative AI tools were used during manuscript preparation to assist with code development and verification of intermediate algebraic steps. All theoretical formulations, results, and physical interpretations were independently developed and validated by the author.
Conflicts of Interest
The author declares no competing financial or non-financial interests.